OpenStudy (anonymous):

A membership to Movie Night Movie Club costs $10 plus $2 per movie. Y = 2x +10. Type how to graph the equation using the y-intercept and slope. Be sure to tell where you place the first point and how to find the location of the second point.

OpenStudy (misssunshinexxoxo):

Your first answer is correct. The y-intercept is 10. The second answer is 20, but I'm going to show you how to find it. When you are asked to find the y-intercept, all you have to do is check and make sure the problem is in y=mx + b format. The problem y= 6x + 20 already is, so you're in luck. Once a problem is in y= mx + b format, it's pretty simple to find the y-intercept because it is always going to be "b" in the equation. Here's a quick example: 5y - 10x = 30 is the equation. Once you get it into y= mx + b format, it is y = 2x + 6. So the y-intercept is 6. Another way to find the y-intercept in a problem, is to substitute 0 in the place of x. So in y= 6x + 20, replace x with 0. You get, y= 20 which is the y-intercept. You can also use this to find the x-intercept, but instead of replacing x with 0, replace y with 0 and solve.

OpenStudy (anonymous):

thank you ! c: